I need help on this homework problem please:
The ratio of trucks to vans at a rental lot was 9:11. If the lot had a total of 80 trucks and vans, how many trucks were on the lot?
thank you in advance.

Answer by richwmiller(17219)

9+11=20
80/20=4
9*4=36 trucks
11*4=44 vans
check
36+44=80